You are given the following data on the balance sheet of a company:

Assets

455,000

Debt

155,000

Equity

300,000

Total assets

455,000

Total liabilities

455,000

The cost of debt capital is 4% and the cost of equity is 13%. The risk free rate is 3% and the expected return on the market index is 7%.

  1. What is the weighted average cost of capital for the company?
  2. What is the beta of the assets of the company?
  3. Suppose the firm changes its borrowing to a debt/equity ratio of 1 from its current level of borrowing. The cost of debt remains at 4%. What is the cost of equity after the change? You should assume that Modigliani-Miller’s irrelevance of capital structure theorem holds.
  4. Discuss what may lead to a breakdown of the Modigliani-Miller irrelevance theorem, and give an overview of the trade-off theory and the pecking-order theory of capital structure.
Answer #1

1.
=155/455*4%+300/455*13%=9.93406593406593%

2.
=(13%-3%)/(7%-3%)*1/(1+155/300)=1.64835164835165

3.
=(13%+4%*155/300)/(1+155/300)+((13%+4%*155/300)/(1+155/300)-4%)*1=15.868%

4.
If transaction costs are present

Trade off theory says that capital strcuture is depednent on the trade off between advantages of interest tax shield versus financial distress costs

Pecking order theory says that there isa hierarchy followed at the time of need of funds